OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. 3y – 2x = 11 y + 2x = 9

OpenStudy (anonymous):

how would i do this?

OpenStudy (compassionate):

You first want to cancel terms, that is, isolate y. Substitution, like elimination, is a method for solving a system of equations. You want to add both the equations together, so that x cancels out. In this equation, we already see +2 and -2 will cancel when added. 3y – 2x = 11 y + 2x = 9 = 4y = 20. You see, the 2s canceled, 3y + y = 4y, and 11 + 9 = 20. Now solve for y by dividing. y = 5. Alright, now take y = 5 and plug it into either equation. I'll choose the second one. y + 2x = 9 5 + 2x = 9 2x = 14 x = 7. So now we know what x equals, and so far we have (7, y) Now plug x into the equation to solve for y. y + 2x = 9 y + 2(7) = 9 y + 14 = 9 y = -5 Thus: (7, -5) Hope this helped ~

OpenStudy (anonymous):

thx sooo much!

OpenStudy (anonymous):

this will help me solve the next 20q :D